OpenStudy (anonymous):

write 2x-4y+8=0 in slope intercept form.

OpenStudy (anonymous):

2x-4y+8=0 Subtract 2x and 8 to the other side -4y=-2x-8 Divide each side by -4 y= 1/2+2 is your answer.